tyswist:

On a scale map, 4 inches represents 60 miles. If Stockton is 150 miles from Logan, how far apart are they on the map?

carmelle:

You can set up this question as a proportion: \[\frac{ 60 }{ 4 }=\frac{ 150 }{ x }\] First, we cross multiply then solve for x. 60 × x 150 × 4 \[60x=600\] \[\frac{ 60x }{ 60 }=\frac{ 600 }{ 60 }\] \[x=10\] This means Stockton is 10 inches away from Logan on a map.